An observer traveling with a constant velocity of 20m/s passes close to a stationary source of sound and noticed that there is a change of frequency of 50Hz as he passes the source....what is the frequency of the source(speed of sound in air=340m/s)


1
Expert's answer
2021-03-04T11:51:44-0500
"f_L=\frac{v+v_L}{v+v_S}f_S\\f_S+50=\frac{340+20}{340+0}f_S\\
f_S=850\ Hz"